solve each proportion
1. n/6= 25/30
2. 6.1/x= 2/3
3. 100/25= t/4
I need help ASAP!
Help!

Answer by stanbon(75887)

solve each proportion
1. n/6= 25/30
Multiply both sides by 6 to get:
n = 6(25/30)
n = (6/30)*25
n = (1/5)*25
n = 5
==============================
2. 6.1/x= 2/3
Invert both sides:
x/6.1 = 3/2
Multiply both sides by 6.1 to get:
x = 6.1(3/2)
x = 18.3/2
x = 9.15
=========================
3. 100/25= t/4
Multiply both sides by 4 to get:
t = 4(100/25)
t = 4*4
t = 16
================
cheers,
Stan H.